# Heat dissipation in marginally stable linear time-delayed Langevin systems

### v3

@@ Introduction
Substituting Eqs. (10) and (11) into Eq. (25), we obtain
$\left\langle\frac{dq}{dt}\right\rangle=aT\cdot\dfrac{1+a\tau}{1-a\tau}\ \ (t\rightarrow\infty)_{.}$ (26)
When calculating Eq. (26), the linearly increasing second moments in Eq. (25) cancel out, resulting in a constant heat dissipation rate in the long-time limit (Fig. 3(a)). As shown in Fig. 3(b), a larger $\tau$ leads to larger heat dissipation rate. The heat dissipation can even become negative for $a<0$ and $\tau<-1/a$, despite the diffusive nature of the dynamics. For $a>0$, the heat dissipation rate diverges as $\tau$ approaches the stability boundary $a\tau=1$.
Figure 3: (a) Time evolution of the average heat dissipation rate in diffusive critical dynamics. Solid line: analytical results (calculated by Eqs. (7), (8) and  (25), where $x_{0}(t)$ is calculated through the numerical integral of Eq. (31) ); dashed line: the asymptotic behavior (Eq. (26)); crosses: numerical simulations with $10^{6}$ trajectories of precision $dt=10^{-4}$, where each point of $\Delta q/\Delta t$ is averaged over interval $\Delta t=0.1$. Here a=5a=-5, $b=5$, $\tau=1$, $T=0.1$, and $\phi(t)=0\ (t\leq 0)$. (b) Average heat dissipation rate of the diffusive critical dynamics in the long-time limit. Here $T=1$.
Figure 4: (a) Time evolution of the average heat dissipation rate in oscillatory critical dynamics. Solid line: the analytical result (calculated by Eqs. (7), (13), and (25), where $x_{0}(t)$ is calculated through the numerical integral of Eq. (31) ); dashed line: the asymptotic behavior (Eq. (27)); crosses: numerical simulations with $10^{6}$ trajectories of precision $dt=10^{-4}$, where each point of $\Delta q/\Delta t$ is averaged over interval $\Delta t=0.1$. Here a=2a=-2, b=2.1b=-2.1, $\tau\approx 4.422$, $T=0.1$, and $\phi(t)=0\ (t\leq 0)$. (b) Divergence speed of the average heat dissipation rate in oscillatory critical dynamics (Eq. (30)). Here b<|a|b<-|a| and $T=1$.
(ii) Oscillatory criticality: Substituting Eqs. (16) and (II) into Eq. (25), we obtain
$\left\langle\frac{dq}{dt}\right\rangle$ $=\frac{4(b^{2}-a^{2})T}{1+b^{2}\tau^{2}-2a\tau}\cdot t\ +A_{0}(t)+B_{0}\ \ (t\rightarrow\infty),$ (27)
@@
$\ \ +(b^{2}-a^{2})C_{1}-aT.$ (29)
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\left\langle\frac{dq}{dt}\right\rangle=\frac{4(b^{2}-a^{2})T}{1+b^{2}\tau^{2}-2a\tau}.$ (30)
The heat dissipation rate of the dynamics of oscillatory criticality linearly diverges in time accompanied by an oscillation, as shown in Fig. 4(a). Its divergence rate (Eq. (30)) increases with $|b|$ for a fixed $a$, and vice versa (Fig. 4(b)). This shows a clear contrast to the heat dissipation of diffusive critical dynamics, although both dynamics show a variance linearly diverging in time.
The sharp difference in heat dissipation between the two types of criticality stems from their underlying dynamical details. The dynamics of oscillatory criticality shows oscillation with a linearly growing amplitude on average. For the majority of trajectories, the motion becomes increasingly violent as the distance covered by oscillation within a period increases with time . The resistance to the noise by such motion ends with linearly divergent heat dissipation in time. On the other hand, the dynamics of diffusive criticality manifests as a scaled Brownian motion. After a sufficiently long time, the range of diffusion remains local within a fixed interval, which leads to bounded dissipation.
@@
Figure 5: Probability distribution of the heat dissipation rate $\Delta q/\Delta t$ for diffusive critical dynamics in the long-time limit, where each sample of $\Delta q/\Delta t$ is averaged over $\Delta t=0.1$ . We take $10^{6}$ samples with simulation precision $dt=10^{-3}$ and set $T=0.1$. (a) Varying $a$ with a=b<0a=-b<0 and $\tau=1$. (b) Varying $a$ with a=b>0a=-b>0 and $\tau=1$. (c) Varying $\tau$ with a=2a=-2 and $b=2$. (d) Varying $\tau$ with $a=0.5$ and b=0.5b=-0.5.
Furthermore, we numerically examine the probability distribution of heat dissipation rate for two types of criticality. The distribution of both criticalities exhibits a single peak and an exponential asymptotic tail. For diffusive criticality, in contrast to the linearly divergent variance of the position, the distribution of heat dissipation converges to an asymptotic one in the long-time limit (Fig. 5). Larger $|a|$ and $|b|$ lead to a broader distribution of heat dissipation (Figs. 5(a) and 5(b)). The distribution also broadens as $\tau$ increases, relaxing to an asymptotic form in the large-$\tau$ limit for a=b<0a=-b<0 (Fig. 5(c)), but spreading without bound as $\tau$ approaches the stability boundary $a\tau=1$ for a=b>0a=-b>0 (Fig. 5(d)). For oscillatory criticality, the distribution continues to spread with time without reaching an asymptotic form (Fig. 6(a)). For a fixed $a$, a larger $|b|$ (and thus a smaller $\tau$ according to the marginal-stability conditions) results in a broader distribution (Figs. 6(b) and 6(c)). These features are consistent with earlier analytical results for $\left\langle dq/dt\right\rangle$, but also show how the variance of heat dissipation rate changes with the relevant parameters.
@@
Figure 6: Probability distribution of the heat dissipation rate $\Delta q/\Delta t$ for the oscillatory critical dynamics at finite time $t$, where each sample of $\Delta q/\Delta t$ is averaged over $\Delta t=0.1$. We take $10^{6}$ samples with simulation precision $dt=10^{-3}$ and set $T=0.1$. (a) Varying the evolution time $t$ with a=2a=-2, b=3b=-3, and $\tau\approx 2.301$. (b) Varying $b$ with a=2a=-2 and $t=70$. (c) Varying $b$ with $a=2$ and $t=70$.

### v4

@@ Introduction
Substituting Eqs. (10) and (11) into Eq. (24), we obtain at large times $t\gg 1/\beta_{1}$:
$\left\langle\frac{dq}{dt}\right\rangle=\frac{aT}{{\color[rgb]{0,0,0}\gamma}}\cdot\dfrac{{\color[rgb]{0,0,0}\gamma}+a\tau}{{\color[rgb]{0,0,0}\gamma}-a\tau}+{\color[rgb]{0,0,0}O(e^{-\beta_{1}t})}.$ (25)
Under the condition $a+b=0$ specified in Eq. (8), the linearly increasing terms in the second moments of Eq. (24) cancel out when calculating Eq. (25). Consequently, the heat dissipation rate becomes asymptotically constant (Fig. 3(a)) and independent of the history at large times $t\gg 1/\beta_{1}$. As shown in Fig. 3(b), a larger $\tau$ leads to larger heat dissipation rate. The heat dissipation can even become negative for $a<0$ and $\tau<-{\color[rgb]{0,0,0}\gamma}/a$, despite the diffusive nature of the dynamics. For $a>0$, the heat dissipation rate diverges as $\tau$ approaches the stability boundary $a\tau={\color[rgb]{0,0,0}\gamma}$. Note that the average heat dissipation vanishes when $\tau=0$. In this specific limit, the cross-term $\langle bx(t-\tau)\circ\xi(t)\rangle$ in Eq. (III), which is zero for $\tau>0$, now yields a finite contribution of $bT$. This additional term is exactly canceled by other components in Eq. (III), resulting in zero dissipation. This unphysical discontinuity in the limit $\tau\rightarrow 0$ (Fig. 3(b)) stems from the idealized assumption of delta-correlated noise in the model.
Figure 3: (a) Time evolution of the average heat dissipation rate in diffusive critical dynamics. Solid line: analytical results [calculated by Eqs. (7), (8) and  (24), where $x_{0}(t)$ is calculated through the numerical integral of Eq. (63)]; dashed line: the asymptotic behavior (Eq. (25)); crosses: numerical simulations with 5000050000 trajectories of precision $dt=10^{-4}$, where $\langle\Delta q/\Delta t\rangle$ is obtained by averaging the values of $\langle dq/dt\rangle$ within a interval $\Delta t=0.1$ centered at each specified time. Red error bars indicate the standard error of the mean. Here a=5a=-5, $b=5$, $\tau=1$, $T=0.1$, $\color[rgb]{0,0,0}\gamma=1$}, and $\phi(t)=0\ (t\leq 0)$. (b) Average heat dissipation rate of the diffusive critical dynamics in the long-time limit. Here $T=1$ and $\color[rgb]{0,0,0}\gamma=1$}.
Figure 4: (a) Time evolution of the average heat dissipation rate in oscillatory critical dynamics. Solid line: the analytical result (calculated by Eqs. (7), (13), and (24), where $x_{0}(t)$ is calculated through the numerical integral of Eq. (63) ); dashed line: the asymptotic behavior (Eq. (26)); crosses: numerical simulations with 5000050000 trajectories of precision $dt=10^{-4}$, where $\langle\Delta q/\Delta t\rangle$ is obtained by averaging the values of $\langle dq/dt\rangle$ within a interval $\Delta t=0.1$ centered at each specified time. Red error bars indicate the standard error of the mean. Here a=2a=-2, b=2.1b=-2.1, $\tau\approx 4.422$, $T=0.1$, $\color[rgb]{0,0,0}\gamma=1$}, and $\phi(t)=0\ (t\leq 0)$. (b) Divergence speed of the average heat dissipation rate in oscillatory critical dynamics (Eq. (29)). Here b<|a|b<-|a|, $T=1$, and $\color[rgb]{0,0,0}\gamma=1$}.
(ii) Oscillatory criticality: We set $\phi(t)=0\ \ (t<0)$ for simplicity, which does not affect the main conclusion. Substituting Eqs. (16) and (II) into Eq. (24), we obtain at large times $t\gg 1/\beta_{1}$:
$\left\langle\frac{dq}{dt}\right\rangle$ $=\frac{4(b^{2}-a^{2})T}{{\color[rgb]{0,0,0}\gamma^{2}}+b^{2}{\color[rgb]{0,0,0}\tau_{c}}^{2}-2{\color[rgb]{0,0,0}\gamma}a{\color[rgb]{0,0,0}\tau_{c}}}\cdot t\ +A_{0}(t)+B_{0}{\color[rgb]{0,0,0}+O(e^{-\beta_{1}t})},$ (26)
@@
$\ \ +\frac{(b^{2}-a^{2})C_{1}}{{\color[rgb]{0,0,0}\gamma}}-\frac{aT}{{\color[rgb]{0,0,0}\gamma}}.$ (28)
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\left\langle\frac{dq}{dt}\right\rangle=\frac{4(b^{2}-a^{2})T}{{\color[rgb]{0,0,0}\gamma^{2}}+b^{2}{\color[rgb]{0,0,0}\tau_{c}}^{2}-2{\color[rgb]{0,0,0}\gamma}a{\color[rgb]{0,0,0}\tau_{c}}}.$ (29)
A general history function $\phi(t)$ does not influence the asymptotic behavior Eq. (29) and only changes the term $A_{0}(t)$ in Eq. (26) through the expression of $\langle x(t)\rangle$. The heat dissipation rate of the dynamics of oscillatory criticality linearly diverges in time accompanied by an oscillation, as shown in Fig. 4(a). Its divergence rate [Eq. (29)] increases with $|b|$ for a fixed $a$ and increases with $a$ for a fixed $|b|$ (Fig. 4(b)). This shows a clear contrast to the heat dissipation of diffusive critical dynamics, although both dynamics show a variance linearly diverging in time.
The sharp difference in heat dissipation between the two types of criticality stems from their underlying dynamical details. The dynamics of oscillatory criticality shows oscillation with a linearly growing amplitude on average. For the majority of trajectories, the motion shows an increasing maximum speed, leading to a greater distance covered per oscillation period over time. The resistance to the noise by such motion ends with linearly divergent heat dissipation in time. On the other hand, the dynamics of diffusive criticality manifests as a scaled Brownian motion. After a sufficiently long time, the range of diffusion remains local within a fixed interval, which leads to bounded dissipation.
@@
Figure 5: Probability distribution of the heat dissipation rate $\Delta q/\Delta t$ for diffusive critical dynamics after a sufficiently long time (at $t=70$), where each sample of $\Delta q/\Delta t$ is averaged over $\Delta t={\color[rgb]{0,0,0}1}$ . We take $10^{6}$ samples with simulation precision $dt=10^{-3}$, and set $T=0.1$ and $\color[rgb]{0,0,0}\gamma=1$}. (a) Varying $a$ with a=b<0a=-b<0 and $\tau=1$. (b) Varying $a$ with a=b>0a=-b>0 and $\tau=1$. (c) Varying $\tau$ with a=2a=-2 and $b=2$. (d) Varying $\tau$ with $a=0.5$ and b=0.5b=-0.5.
Furthermore, we numerically examine the probability distribution of heat dissipation rate for two types of criticality. The distribution of both criticalities exhibits a single peak and an exponential asymptotic tail. For diffusive criticality, in contrast to the linearly divergent variance of the position, the distribution of heat dissipation converges to an asymptotic one after a sufficiently long time (Fig. 5). Larger $|a|$ and $|b|$ lead to a broader distribution of heat dissipation (Figs. 5(a) and 5(b)). The distribution also broadens as $\tau$ increases, relaxing to an asymptotic form in the large-$\tau$ limit for a=b<0a=-b<0 (Fig. 5(c)), but spreading without bound as $\tau$ approaches the stability boundary $a\tau={\color[rgb]{0,0,0}\gamma}$ for a=b>0a=-b>0 (Fig. 5(d)). For oscillatory criticality, the distribution continues to spread with time without reaching an asymptotic form (Fig. 6(a)). For a fixed $a$, a larger $|b|$ (and thus a smaller $\tau$ according to the marginal-stability conditions) results in a broader distribution (Figs. 6(b) and 6(c)). These features are consistent with earlier analytical results for $\left\langle dq/dt\right\rangle$, but also show how the variance of heat dissipation rate changes with the relevant parameters.
@@
Figure 6: Probability distribution of the heat dissipation rate $\Delta q/\Delta t$ for the oscillatory critical dynamics at finite time $t$, where each sample of $\Delta q/\Delta t$ is averaged over $\Delta t=1$. We take $10^{6}$ samples with simulation precision $dt=10^{-3}$, and set $T=0.1$ and $\color[rgb]{0,0,0}\gamma=1$}. (a) Varying the evolution time $t$ with a=2a=-2, b=3b=-3, and $\tau\approx 2.301$. (b) Varying $b$ with a=2a=-2 and $t=70$. (c) Varying $b$ with $a=2$ and $t=70$.
Asymptotic behaviors of heat dissipation rate near the stability boundaries
In this section, we investigate the asymptotic behaviors of the heat dissipation rate as the dynamics [Eq. (1)] approaches the two classes of critical dynamics from the stable regime. According to Ref. [23], the dynamics in Eq. (1) is stable if and only if either
$a+b<0,\ \ a-b\leq 0,$ (30)
for any $\tau>0$ holds, or
a+b<0,ab>0, $a-b>0,\ \ 0<\tau<\frac{\gamma\arccos(-a/b)}{(b^{2}-a^{2})^{1/2}}$ (31)
is satisfied. In the stable regime, all characteristic roots in Eq. (5) have a negative real part [23]. Therefore, from Eqs. (II), (6) and (7), the system asymptotically reaches a steady state in the long-time limit, in which the solution Eq. (II) follows a history-independent stationary Gaussian distribution with zero mean and finite variance. The average heat dissipation rate of the system [Eq. (1)] in the steady state can be expressed as [31] (see also [26, 27]):
$\langle J\rangle=\frac{aT}{\gamma}+\frac{2T}{\gamma^{2}}[(a^{2}+b^{2})K(0)+2abK(\tau)],$ (32)
where $K(t)\equiv\int_{0}^{\infty}x_{0}(s)x_{0}(s+t)ds$. The explicit expressions for $K(0)$ and $K(\tau)$ are [23]:
[…]
with $l=\sqrt{|a^{2}-b^{2}|}$. In the following, we focus on the behaviors of $\langle J\rangle$ in Eq. (32) near the two critical boundaries defined by Eqs. (8) and (13).
First, we fix $a$ and $\tau$, and vary $b$ towards b=ab=-a. We take $b=-a-\epsilon_{0}$, where $|\epsilon_{0}|\ll|a|$ and according to Eqs. (30) and (31), $\epsilon_{0}>0$. From the expressions of $K(0)$ and $K(\tau)$ in Eqs. (34) and (33), we obtain
$K(0)$ $=-\frac{\gamma^{2}}{2\epsilon_{0}(a\tau-\gamma)}+O(1),\ \$(35)
$K(\tau)$ $=K(0)+\frac{a\tau^{2}}{2(a\tau-\gamma)}-\frac{\tau}{2}+O(\epsilon_{0}).$ (36)
Therefore, using the expression for heat dissipation rate in the stable dynamics:
$\langle J\rangle$ $=\frac{aT}{\gamma}+\frac{2T}{\gamma^{2}}[(a^{2}+b^{2})K(0)+2abK(\tau)]$
$=\frac{aT}{\gamma}-\frac{4a^{2}T}{\gamma^{2}}\left(\frac{a\tau^{2}}{2(a\tau-\gamma)}-\frac{\tau}{2}\right)+O(\epsilon_{0})$
$=\frac{aT}{\gamma}\frac{\gamma+a\tau}{\gamma-a\tau}+O(\epsilon_{0}),$ (37)
which is consistent with the heat dissipation rate for diffusive criticality [Eq. (25)]. On the other hand, for oscillatory criticality, we fix $a$ and $b$ and vary $\tau$ towards $\tau_{c}=\gamma\arccos(-a/b)/(b^{2}-a^{2})^{1/2}$. Let $\tau=\tau_{c}-\epsilon_{1}\ \ (0<\epsilon_{1}\ll\tau_{c})$; then we have
$K(0)=\frac{\gamma^{2}}{(b^{2}-a^{2})\epsilon_{1}}+O(1),\ \ K(\tau)=-\frac{a}{b}K(0)+O(1),$ (38)
and
$\langle J\rangle$ $=\frac{T}{\gamma}a+\frac{2T}{\gamma^{2}}\left[(a^{2}+b^{2})K(0)+2ab\left(-\frac{a}{b}K(0)\right)\right]+O(1)$
$=\frac{2T}{\epsilon_{1}}+O(1),$ (39)
which shows that the heat dissipation rate in the steady state diverges as the dynamics approaches oscillatory criticality. Note that when the parameters cross the stability boundaries from the stable regime, at least one characteristic root acquires a positive real part. As a result, both the dynamical quantities in Eqs. (6) and (7) and the average heat dissipation rate in Eq. (24) grow exponentially with time.
To further understand the asymptotic behaviors near criticality, we examine how the dynamical quantities and the average heat dissipation rate respond to infinitesimal shifts of the characteristic root for the two criticality classes. From the characteristic equation (5), we set
$F(s)=\gamma s-a-be^{-s\tau}.$ (40)
The marginally stable cases correspond to $F(s_{c})=0$, where $s_{c}$ is the characteristic root with a zero real part. In the case of diffusive criticality, we have $s_{c}=0$. We consider the dynamics where $b=-a-\epsilon_{0}\ \ (\epsilon_{0}>0)$ and a perturbed characteristic root $s_{0}=s_{c}+\Delta s_{0}$. To first order, we obtain:
$\left.\frac{\partial F}{\partial s}\right|_{(s=0,\ b=-a)}\cdot\Delta s_{0}-\left.\frac{\partial F}{\partial b}\right|_{(s=0,\ b=-a)}\cdot\epsilon_{0}\approx 0,$ (41)
which yields
$\Delta s_{0}\approx-R_{0}\epsilon_{0}=O(\epsilon_{0})<0,$ (42)
where
$R_{0}\equiv\frac{1}{\gamma-a\tau}.$ (43)
While for oscillatory criticality, we have $s_{c}=\pm i\omega_{c}$. Without loss of generality, let $s_{c}=i\omega_{c}$. For the dynamics with $\tau=\tau_{c}-\epsilon_{1}\ (\epsilon_{1}>0)$, we consider a perturbed characteristic root $s_{1}=s_{c}+\Delta s_{1}$ and define $\Delta s_{1}=\mathrm{Re}[s_{1}]+i(\omega_{2}-\omega_{c})$. We then obtain to first order:
$\left.\frac{\partial F}{\partial s}\right|_{(i\omega_{c},\tau_{c})}\cdot\Delta s_{1}-\left.\frac{\partial F}{\partial\tau}\right|_{(i\omega_{c},\tau_{c})}\cdot\epsilon_{1}\approx 0,$ (44)
which leads to
$\Delta s_{1}$ $\approx\frac{\left(\frac{\partial F}{\partial\tau}\right)\epsilon_{1}}{\frac{\partial F}{\partial s}}=\frac{bi\omega_{c}e^{-i\omega_{c}\tau_{c}}\cdot\epsilon_{1}}{\gamma+b\tau_{c}e^{-i\omega_{c}\tau_{c}}}\$
$=\frac{(-\gamma\omega_{c}^{2}-ia\omega_{c})\epsilon_{1}}{(\gamma-a\tau_{c})+i\gamma\omega_{c}\tau_{c}}.$ (45)
In the last line, we have used $be^{-i\omega_{c}\tau_{c}}=i\gamma\omega_{c}-a$, which follows from $F(s=i\omega_{c},\ \tau=\tau_{c})=0$. This shows
$\mathrm{Re}[s_{1}]\approx-R_{1}\epsilon_{1}=O(\epsilon_{1})<0,$ (46)
where
$R_{1}\equiv\frac{\gamma^{2}\omega_{c}^{2}}{(\gamma-a\tau_{c})^{2}+\gamma^{2}\omega_{c}^{2}\tau_{c}^{2}}.$ (47)
Therefore, for stable dynamics close to either type of criticality, it follows from Eqs. (6), (7) and (101) that
$\frac{d}{dt}\langle x^{2}(t)\rangle=\frac{2T}{\gamma}x_{0}^{2}(t)=O(e^{-2\varepsilon t})+O(e^{-(\varepsilon+|\mathrm{Re}[\nu]|)t})+\cdots,$ (48)
where $\varepsilon=R_{0}\epsilon_{0}$ (or $\varepsilon=R_{1}\epsilon_{1}$), and $\nu=\nu_{0}$ (or $\nu=\nu_{1}$) is the characteristic root with the second largest real part for the dynamics close to diffusive (or oscillatory) criticality. This shows that in the limit $\epsilon_{0}\rightarrow 0$ or $\epsilon_{1}\rightarrow 0$, the dynamics close to either diffusive or oscillatory criticality requires asymptotically infinite time (of order $O(1/\epsilon_{0})$ or $O(1/\epsilon_{1})$, respectively) to achieve the steady state. However, the rate of convergence for the average heat dissipation rate differs between the two classes of criticality. From Eqs. (24) and (48), for the dynamics close to diffusive criticality, we have
$\frac{d}{dt}\left\langle\frac{dq(t)}{dt}\right\rangle=\frac{1}{\gamma}\frac{d}{dt}\bigg(b^{2}[\langle x^{2}(t-\tau)\rangle-\langle x^{2}(t)\rangle]$
$\quad\quad\quad+(b^{2}-a^{2})\langle x^{2}(t)\rangle+a\gamma\frac{d\langle x^{2}(t)\rangle}{dt}-aT\bigg)$
$=O(R_{0}\epsilon_{0}e^{-2R_{0}\epsilon_{0}t})+O([R_{0}\epsilon_{0}+|\nu_{0}|]e^{-(R_{0}\epsilon_{0}+|\mathrm{Re}[\nu_{0}]|)t})+\cdots.$ (49)
In the last line of Eq. (IV), we have used
$b^{2}[\langle x^{2}(t-\tau)\rangle-\langle x^{2}(t)\rangle]=\frac{2Tb^{2}}{\gamma}\int_{t-\tau}^{t}x_{0}^{2}(s)ds$
$\quad=O(b^{2}\tau\cdot e^{-2R_{0}\epsilon_{0}t})+\cdots,$ (50)
$b^{2}-a^{2}=2a\epsilon_{0}+O(\epsilon_{0}^{2}).$ (51)
On the other hand, for the dynamics close to the oscillatory criticality, we have
$\frac{d}{dt}\left\langle\frac{dq(t)}{dt}\right\rangle=O(e^{-2R_{1}\epsilon_{1}t})+O(e^{-(R_{1}\epsilon_{1}+|\mathrm{Re}[\nu_{1}]|)t})+\cdots,$ (52)
because in this case $b^{2}-a^{2}\sim O(1)$ and $(b^{2}-a^{2})\cdot d\langle x^{2}(t)\rangle/dt$ accounts for the leading-order contribution in Eq. (52). Therefore, in the limit $\epsilon_{0}\rightarrow 0$ or $\epsilon_{1}\rightarrow 0$, the heat dissipation rate of stable dynamics near diffusive criticality approaches a constant on a finite time scale of order $O(1/|\mathrm{Re}[\nu_{0}]|)$, whereas that near oscillatory criticality requires an asymptotically infinite time of order $O(1/\epsilon_{1})$ to reach the specific steady-state value given by the expression Eq. (32) for stable dynamics.
We also investigate the spectral decomposition of the steady-state heat dissipation rate as the parameters asymptotically approaches the two critical boundaries. According to the Harada-Sasa equality, the steady-state heat dissipation rate can be decomposed as follows [54]:
$\left\langle J\right\rangle=\gamma\left\{v_{s}^{2}+\int_{-\infty}^{\infty}\left[\tilde{C}_{v}(\omega)-2T\tilde{R}^{\prime}_{v}(\omega)\right]\frac{d\omega}{2\pi}\right\}.$ (53)
Here, $v_{s}$ is the steady-state velocity, $\tilde{C}_{v}(\omega)$ is the velocity power spectral density and $\tilde{R}_{v}^{\prime}(\omega)$ is the real part of linear response function. The integrand $\tilde{C}_{v}(\omega)-2T\tilde{R}^{\prime}_{v}(\omega)$ quantifies the violation of the fluctuation-response relation, which vanishes for all $\omega$ in equilibrium systems but is generally non-zero in non-equilibrium steady states [55, 56]. The Harada-Sasa equality has been shown to remain valid for time-delayed Langevin systems [57, 31].
(a) a=2a=-2
(b) a=2a=-2
(c) $a=0.7$
Figure 7: Spectral decomposition of the heat dissipation rate in the steady state in linear time-delayed Langevin equation [Eq. (1)] near the diffusive criticality. Note that (a) and (b) show the same curves with different range of $\omega$. Here, we fix $\tau=1$, $T=0.1$, and $\gamma=1$.
In particular, for the linear time-delayed dynamics [Eq. (1)] under the stability condition Eq. (30) or (31), the steady-state velocity is $v_{s}=0$. The integrand in the spectral decomposition of the steady-state heat dissipation rate is given by [31]:
$\tilde{C}_{v}(\omega)-2T\tilde{R}^{\prime}_{v}(\omega)$
$\quad=-2T\frac{b\omega\sin(\omega\tau)}{[a+b\cos(\omega\tau)]^{2}+[\gamma\omega+b\sin(\omega\tau)]^{2}}.$ (54)
As shown in Fig. 7, the shape of the spectrum converges to a limiting form when $\epsilon_{0}\rightarrow 0$ near the diffusive criticality. Near $\omega=0$, the spectrum exhibits a steep rise (or drop) with an asymptotically divergent slope for $a>0$ (or $a<0$), as seen in Figs. 7(b) and 7(c). Expanding Eq. (54) for $\omega\ll 1/\tau$ to first order yields
$\tilde{C}_{v}(\omega)-2T\tilde{R}^{\prime}_{v}(\omega)\approx\frac{-2Tb\tau\omega^{2}}{\epsilon_{0}^{2}+(\gamma+b\tau)^{2}\omega^{2}}.$ (55)
Thus, for $\omega\ll\epsilon_{0}/|\gamma+b\tau|$, we have $\tilde{C}_{v}(\omega)-2T\tilde{R}^{\prime}_{v}(\omega)\approx-2Tb\tau\omega^{2}/\epsilon_{0}^{2}$. This indicates a sharp change from zero near $\omega=0$, with the spectrum reaching an extremum of approximately $2Tb\tau/(\gamma+b\tau)^{2}$ when $\epsilon_{0}/|\gamma+b\tau|\ll\omega\ll 1/\tau$. The value of $\epsilon_{0}$ has a limited effect on the spectrum for $\omega\gg 1/\tau$. In contrast, for parameters near the oscillatory criticality, a resonance peak emerges near $\omega=\omega_{c}=\sqrt{b^{2}-a^{2}}/\gamma$ in the spectrum. As shown in Fig. 8, the peak height increases as $\tau$ approaches $\tau_{c}=\gamma\arccos(-a/b)/(b^{2}-a^{2})^{1/2}$ from below. To analyze this, we set $\tau=\tau_{c}-\epsilon_{1}$ with $0<\epsilon_{1}\ll\tau_{c}$ and examine the spectrum at $\omega=\omega_{c}+\mu$, where $|\mu|\ll\omega_{c}$. Analytically, expanding $\omega\tau$ to first order, the spectrum near the resonance peak at $\omega=\omega_{c}+\mu$ can be approximated as
$\tilde{C}_{v}(\omega)-2T\tilde{R}^{\prime}_{v}(\omega)\approx-\frac{2Tb\omega_{c}\sin(\omega_{c}\tau_{c})}{D_{1}\mu^{2}+D_{2}\mu\epsilon_{1}+D_{3}\epsilon_{1}^{2}},$
$=\frac{2T(b^{2}-a^{2})}{\gamma[D_{1}(\mu-\mu_{\mathrm{max}})^{2}+(b^{2}-a^{2})^{2}\epsilon_{1}^{2}/D_{1}]},$ (56)
where
$D_{1}$ $\equiv\gamma^{2}\omega_{c}^{2}\tau_{c}^{2}+(\gamma+a\tau_{c})^{2},$ (57)
$D_{2}$ $\equiv-2\omega_{c}(b^{2}\tau_{c}+a\gamma),$ (58)
$D_{3}$ $\equiv b^{2}\omega_{c}^{2},$ (59)
$\mu_{\mathrm{max}}$ $=-\frac{D_{2}\epsilon_{1}}{2D_{1}}.$ (60)
The resonance peak has a Lorentzian line shape. Its maximum value is
$\tilde{C}_{v}(\omega)-2T\tilde{R}^{\prime}_{v}(\omega)\bigg|_{\mu=\mu_{\mathrm{max}}}=\frac{2TD_{1}}{\gamma(b^{2}-a^{2})\epsilon_{1}^{2}}>0,$ (61)
and its full width at half maximum is
$\Delta\mu=\frac{2(b^{2}-a^{2})\epsilon_{1}}{D_{1}}.$ (62)
These results show that as $\epsilon_{1}\rightarrow 0$, the resonance peak becomes increasingly narrow and tall, with its integrated area diverging asymptotically. By combining Eqs. (61) and (62) and integrating the Lorentzian curve using Eq. (53), we confirm that the asymptotically divergent term $2T/\epsilon_{1}$ in Eq. (IV) originates precisely from the integrated area of this resonance peak.
(a) a=2a=-2
(b) a=2a=-2
Figure 8: Spectral decomposition of the heat dissipation rate in the steady state in linear time-delayed Langevin equation [Eq. (1)] near the oscillatory criticality. Here, we fix a=2a=-2, b=3b=-3, $T=0.1$ and $\gamma=1$. Note that in these cases, $\tau_{c}\approx 1.0288$ and $\omega_{c}\approx 2.236$.